Sir , jee mains 2 nd shift april 12 question from atoms and nuclei , what is its answer? Can u plz help me out?

Grade:12th pass

1 Answers

sree lakshmi
4 years ago

Electron in hydrogen atom first jumps from third excited state to second excited state and then from second excited to the first excited state. The ratio of the wavelengths λ1:λ2 emitted in the two cases is 

ans; 1st the electron jumps from 3rd excited state to the 2nd ie from n=4 to n=3 
then 1/λ1 =R(1/n12 -1/n22 )= R ( 1/9 – 1/16) =R (7/144)
similarly 1/λ2 = R ( ¼ – 1/9 )=R(5/36)
∴ λ1/λ2 = 1/R(144/7) * R (5/36) =4*5/7 =20/7
so the ans is 20/7